OpenStudy (anonymous):

Evaluate the following expression for x=–2 and y=–4. 4(x-y)^4

OpenStudy (anonymous):

A. –64 B. 64 C. –5,184 D. 5,184

OpenStudy (oaktree):

So you just want to plug in your x and y. You have\[4(-2-(-4))^4 = 4(-2+4)^4 = 4(2)^4 = 4*16 = 64\]Get it?

OpenStudy (anonymous):

Yes, thank you.